# This is a preprint submitted to: Archives of Computational Methods in Engineering

### v1

@@ Abstract
The rapidly evolving field of engineering design of functional surfaces necessitates sophisticated tools to manage the inherent complexity of high-dimensional design spaces. This review delves into the field of design-space dimensionality reduction techniques tailored for shape optimization, bridging traditional methods and cutting-edge technologies. Dissecting the spectrum of these techniques, from classical linear approaches like principal component analysis to more nuanced nonlinear methods such as autoencoders, the discussion extends to innovative physics-informed methods that integrate physical data into the dimensionality reduction process, enhancing the predictive accuracy and relevance of reduced models. By integrating these methods into optimization frameworks, it is shown how they significantly mitigate the curse of dimensionality, streamline computational processes, and refine the exploration and optimization of complex functional surfaces. The survey provides a classification of method and highlights the transformative impact of these techniques in simplifying design challenges, thereby fostering more efficient and effective engineering solutions.
Keywords Shape optimization  $\cdot$ Space reduction  $\cdot$ Dimensionality reduction  $\cdot$ Representation learning  $\cdot$ Principal component analysis  $\cdot$ Parametric model embedding  $\cdot$ Active subspace  $\cdot$ Autoencoders
@@ 1 Introduction
Shape optimization is integral to advancing performance and efficiency in a myriad of engineering disciplines, including aerodynamics, hydrodynamics, heat transfer, and turbomachinery. The optimization process is crucial not only for enhancing functional outputs but also for meeting increasingly stringent environmental regulations and safety standards. Each of these fields involves navigating a complex, high-dimensional design space defined by an extensive array of design variables and constraints, which complicates the optimization process and amplifies computational demands.
@@
The challenge is further intensified by the so-called curse of dimensionality [1], which refers to the exponential growth in computational resources required as the number of dimensions (variables) in the system increases. To manage this complexity, dimensionality reduction techniques have emerged as a cornerstone strategy. By effectively reducing the number of variables under consideration without significant loss of critical information, these techniques simplify the design space, making it more tractable for simulation and analysis.
@@
This survey paper provides a comprehensive review of dimensionality reduction techniques applied in the context of shape optimization. It provides a first classification of methods and explores a variety of approaches from traditional linear to more advanced nonlinear methods. Each technique’s theoretical foundations, applications, and impacts on the optimization process are critically examined, highlighting their integration within simulation-based design optimization frameworks.
@@
The aim is to delineate the roles these techniques play in different engineering applications, emphasizing how they can be harnessed to streamline computational processes, enhance accuracy, and expedite the exploration of optimal designs. Through detailed discussions and case studies, this paper seeks to equip practitioners and researchers with the knowledge to effectively implement dimensionality reduction techniques in their respective fields, fostering innovation and efficiency in engineering design [2, 3].
@@ 2 Shape Optimization Problem Formulation
Consider a manifold $\mathcal{G}$, which identifies the original/parent shape, whose coordinates in the 3D-space are represented by ${\mathbf{g}}({\boldsymbol{\xi}})\in\mathbb{R}^{3}$; ${\boldsymbol{\xi}}\in\mathcal{G}$ are curvilinear coordinates defined on $\mathcal{G}$. Assume that, for the purpose of shape optimization, ${\mathbf{g}}$ can be transformed to a deformed shape/geometry ${\mathbf{g}}^{\prime}({\boldsymbol{\xi}},{\mathbf{u}})$ by
@@
where ${\boldsymbol{\delta}}({\boldsymbol{\xi}},{\mathbf{u}})\in\mathbb{R}^{3}$ is the resulting shape modification vector, defined by arbitrary shape parameterization or modification method based on fully-parametric or partially-parametric models [4], and ${\mathbf{u}}\in\mathcal{U}\subset\mathbb{R}^{M}$ is the design variable vector, as depicted in Fig. 1.
Figure 1: Shape modification example
@@
$\min_{{\mathbf{u}}}\quad$ $f({\mathbf{u}},{\mathbf{y}})$ (2)
subject to $g_{i}({\mathbf{u}},{\mathbf{y}})\leq 0,\quad i=1,\ldots,I$
and to $h_{j}({\mathbf{u}},{\mathbf{y}})=0,\quad j=1,\ldots,J$
@@
where $f$ is the objective function, ${\mathbf{u}}_{l}$ and ${\mathbf{u}}_{u}$ are the lower and upper bounds for the design variable vector, ${\mathbf{y}}$ are the environmental and/or operational conditions, $g_{i}$ the inequality constraints, and $h_{j}$ the equality constraints. The associated workflow based on the extended design structure matrix (XDSM) [5] is shown in Fig. 2.
Figure 2: Standard simulation-based design optimization XDSM diagram for shape optimization
@@ 3 Tackling the Curse of Dimensionality
The curse of dimensionality, a term coined by Richard Bellman in 1961, describes the exponential increase in complexity that accompanies adding dimensions to a mathematical space [1]. In the context of shape optimization, as the dimensionality $M$ of the design variable vector $\mathbf{u}$ and the design space bounds increase, the volume of the space grows exponentially, making it increasingly difficult to cover it adequately with a finite number of observations. This complexity not only impacts the efficacy of optimization methods, see, e.g., Fig. 3, including those relying on surrogate models and multi-fidelity approaches but also complicates surrogate training and uncertainty quantification.
@@
Figure 4 provides a systematic classification of dimensionality reduction techniques applied in vehicle design optimization. Addressing the curse of dimensionality in such a complex domain requires distinguishing between two principal strategies: space reduction and dimensionality reduction.
@@
Furthermore, design-space dimensionality reduction techniques can be categorized based on their reliance on simulation or geometric data. Simulation-driven techniques, such as indirect methods and some forms of GTM, derive their effectiveness from detailed understanding and manipulation of the objective functions in simulations. Conversely, methods like PCA and its variants primarily utilize geometric data. However, modern approaches often blend these data types, leveraging both geometric and physical simulation data to enhance the applicability and effectiveness of the dimensionality reduction.
The choice of dimensionality reduction technique in shape optimization significantly influences the efficiency and success of the optimization process. By classifying these methods into distinct categories and understanding their underlying principles and applications, engineers and researchers can better select and apply the most appropriate techniques for their specific design challenges. This classification not only aids in comprehending the vast landscape of dimensionality reduction but also highlights the innovative ways these techniques are integrated and applied in the complex field of shape optimization of functional surfaces.
@@ 3.1 Literature overview
To empirically assess the prevalence and efficacy of design space dimensionality reduction methods, a comprehensive bibliographic review was conducted, focusing on their applications across various domains of shape optimization. Non-peer-reviewed articles, such as conference abstracts/papers and editorials, were excluded to ensure the review’s academic rigor. The statistical analysis, supported by visual trends from the literature, reveals distinct patterns in the adoption of direct versus indirect methods, the preference for linear over nonlinear approaches, and the allocation between space reduction and dimensionality reduction techniques. This analysis not only reflects the current state-of-the-art but also guides future research directions by identifying areas lacking in-depth exploration and potential methodological synergies.
@@
When distinguishing between space reduction and dimensionality reduction, as shown in Fig. 6, 78% of the studies focus on dimensionality reduction while only 22% address space reduction. This highlights a predominant trend towards techniques that reduce the number of variables directly, rather than just condensing the space within which the variables exist.
Considering dimensionality reduction methods, the examination of literature trends underscores a significant leaning towards direct dimensionality reduction methods, with a staggering 96% of reviewed articles opting for direct techniques as opposed to only 4% utilizing indirect methods. This suggests a strong preference for methods that actively transform the input data space to achieve dimensionality reduction (see Fig. 6).
@@
In terms of methodological choices for direct dimensionality reduction approaches (see Fig. 8), there is an overwhelming preference for linear techniques, with 87% of the cases applying linear methods over 13% for nonlinear approaches. This could indicate that the simplicity and computational efficiency of linear methods, such as PCA and SVD, continue to make them popular choices in many engineering applications despite the increasing capability of nonlinear methods to capture more complex data relationships.
Finally, analyzing the approaches based on their reliance on data types, as shown in Fig. 8, 50% of the methods are predominantly geometry-based, 43% are simulation-based, and a smaller segment of 7% combines both approaches. This distribution emphasizes the critical role of geometric considerations in the design processes and the significant integration of simulation data to inform these designs.
The temporal trend across the years (see Fig. 9) shows a steady increase in the publications related to dimensionality reduction in shape optimization, with a clear linear upward trend (note that the data for the year 2024 is partial, as the bibliographic research was conducted on March 1, 2024). This growing interest underlines the expanding recognition of the importance of advanced dimensionality reduction techniques in tackling the complexity of modern engineering design problems.
@@ 4 Space Reduction Methods
Space reduction methods play a crucial role in the field of shape optimization, particularly when dealing with high-dimensional design spaces that challenge traditional optimization techniques. These methods enhance the efficiency and effectiveness of optimization processes by strategically narrowing the focus of the design space. This is achieved either through bounds narrowing, which reduces the breadth of the design space to concentrate on regions most likely to contain optimal solutions, or through geometric filtering, which excludes non-viable or abnormal regions based on sophisticated data-driven models or constraints (see Fig. 10). Both approaches aim to refine the search area, ensuring that optimization efforts are concentrated on the most promising or feasible parts of the design space. By focusing on these key strategies, the following sections will explore how bounds narrowing and geometric filtering serve as effective tools in simplifying complex optimization tasks, ultimately leading to more innovative and practical engineering solutions.
@@ 4.1 Bounds narrowing
These methods are instrumental in narrowing the bounds of the design space, traditionally expansive, to focus on smaller subspaces where the optimal design is likely found.
In the optimization of an air-launched satellite launch vehicle [12], bounds narrowing is utilized to focus on the most promising regions within the existing design space. This strategy combines genetic algorithms and simulated annealing, enabling efficient exploration and optimization of complex multidisciplinary design parameters. Similarly, in the context of ship design, space reduction is implemented through set-based approaches [13]. This method differs from traditional optimizations by using sets of design variables, which maintain flexibility and adapt to evolving design requirements, thus facilitating informed decision-making and enhancing the ability to manage complex constraints. Additionally, bounds narrowing has been applied to wing weight minimization [14] and multidisciplinary design problems involving aero-structural decomposition [15]. This approach identifies and explores promising design space regions, significantly reducing the space bounds and refining the search process to enhance computational efficiency and focus optimization efforts where most beneficial. In developing a variable-accuracy global metamodel-based optimization framework [16], bounds narrowing focuses on significant design regions and adapts model fidelity, effectively managing design experiments and uncertainty quantification. This strategy reduces computational complexity and is particularly beneficial for designs based on fluid-structure interaction problems. Aerodynamic shape optimization also benefits from bounds narrowing, where adaptive response surface methodologies and intelligent space exploration strategies [17, 18, 19] target the optimization of airfoil designs. By constructing and refining Kriging-based surrogate models within a reduced design space, this method concentrates on key performance areas, streamlining the exploration and optimization process. In the hydrodynamic shape optimization of a blended-wing-body underwater glider, bounds narrowing is used to focus optimization efforts on critical regions, demonstrating substantial improvements in hydrodynamic performance with enhanced computational efficiency [20]. This approach optimizes sectional airfoils for maximum lift-to-drag ratio, illustrating the method’s effectiveness in navigating complex design spaces. Finally, a novel optimization method combining particle swarm optimization with chaotic logistic maps and a comprehensive learning strategy is proposed in [21]. This method efficiently addresses space reduction by structuring the optimization process into explorative and exploitative subpopulations, which manage the balance between global and local searches effectively. It enhances the ability to find optimal solutions in size and shape optimization tasks for structures by avoiding local optima and refining searches in dynamically reduced design spaces, thereby improving the search process for practical-scale engineering problems.

### v2

@@ Abstract
The rapidly evolving field of engineering design of functional surfaces necessitates sophisticated tools to manage the inherent complexity of high-dimensional design spaces. This survey paper offers a scoping review, i.e., a literature mapping synthesis borrowed from clinical medicine, delving into the field of design-space dimensionality reduction techniques tailored for shape optimization, bridging traditional methods and cutting-edge technologies. Dissecting the spectrum of these techniques, from classical linear approaches like principal component analysis to more nuanced nonlinear methods such as autoencoders, the discussion extends to innovative physics-informed methods that integrate physical data into the dimensionality reduction process, enhancing the physical relevance and effectiveness of reduced design spaces. By integrating these methods into optimization frameworks, it is shown how they significantly mitigate the curse of dimensionality, streamline computational processes, and refine the design exploration and optimization of complex functional surfaces. The survey provides a classification of method and highlights the transformative impact of these techniques in simplifying design challenges, thereby fostering more efficient and effective engineering solutions.
Keywords Shape optimization  $\cdot$ Space reduction  $\cdot$ Dimensionality reduction  $\cdot$ Representation learning  $\cdot$ Principal component analysis  $\cdot$ Parametric model embedding  $\cdot$ Active subspace  $\cdot$ Autoencoders  $\cdot$ Scoping review
@@ 1 Introduction
Shape optimization is integral to advancing performance and efficiency in a myriad of engineering disciplines, including aerodynamics, hydrodynamics, and heat transfer, with applications ranging from turbomachinery to aircraft design, from ships and underwater vehicles to cars and motorcycle, from energy harvesting devices to space rockets. The optimization process is crucial not only for enhancing functional outputs but also for meeting increasingly stringent environmental regulations and safety standards. Each of these fields involves navigating a complex, high-dimensional design space defined by an extensive array of design variables and constraints, which complicates the optimization process and amplifies computational demands.
@@
The challenge is further intensified by the so-called curse of dimensionality [1], which refers to the exponential growth in computational resources required as the number of dimensions (variables) in the problem increases. To manage this complexity, dimensionality reduction techniques have emerged as a cornerstone strategy. By effectively reducing the number of variables under consideration without significant loss of critical information, these techniques simplify the design space, making it more tractable for optimization.
Although dimensionality reduction methods are widely used in various engineering applications, shape optimization problems pose unique challenges and requirements. Unlike typical functional-space applications (e.g., flow fields, stress/strain fields), shape optimization inherently involves geometric design spaces, requiring explicit and careful adaptations of these techniques to address geometric constraints, ensure feasible physical back-projections, and embed meaningful geometric metrics. Thus, a critical analysis specifically tailored to the geometric peculiarities and demands of shape optimization is essential, as traditional dimension reduction methods cannot be applied directly without significant assumptions and methodological modifications.
@@
This survey paper provides a comprehensive scoping review on dimensionality reduction techniques applied in the context of shape optimization. It provides a first classification of methods and explores a variety of approaches from traditional linear to more advanced nonlinear methods. Each technique’s theoretical foundations, applications, and impacts on the shape optimization process are critically examined, highlighting their integration within simulation-based design optimization frameworks.
It is worth clarifying here that, while the term shape optimization broadly encompasses a variety of problems, including structural, topological, and multi-patch manifold optimization, this survey specifically addresses the optimization of external shapes. By definition, these external shapes represent the primary geometric interfaces interacting directly with external physical phenomena, such as fluid dynamics and heat transfer, and their geometric variations strongly and directly influence performance metrics. Structural and topological optimization problems, where geometry-performance relationships may be indirect or less explicit, constitute a distinct class that is not the main focus of this paper, although selected examples of these methods are briefly referenced in the concluding section.
@@
The aim here is to delineate the roles design-space dimensionality reduction techniques for external shape optimization play in different engineering applications, emphasizing how they can be harnessed to streamline computational processes, enhance accuracy, and expedite the exploration of optimal designs. Through detailed discussions and case studies, this paper seeks to equip practitioners and researchers with the knowledge to effectively implement dimensionality reduction techniques for shape optimization in their respective fields, fostering innovation and efficiency in engineering design [2, 3].
@@ 2 Shape Optimization Problem Formulation
This section underlines the formulation of the shape optimization problem discussed in the present survey. Specifically, consider a manifold $\mathcal{G}$, which identifies the original/parent shape, whose coordinates in the 3D-space are represented by ${\mathbf{g}}({\boldsymbol{\xi}})\in\mathbb{R}^{3}$; ${\boldsymbol{\xi}}\in\mathcal{G}$ are curvilinear coordinates defined on $\mathcal{G}$. Assume that, for the purpose of shape optimization, ${\mathbf{g}}$ can be transformed to a deformed shape/geometry ${\mathbf{g}}^{\prime}({\boldsymbol{\xi}},{\mathbf{u}})$ by
@@
where ${\boldsymbol{\delta}}({\boldsymbol{\xi}},{\mathbf{u}})\in\mathbb{R}^{3}$ is the resulting shape modification vector, defined by arbitrary shape parameterization or modification method based on fully-parametric or partially-parametric models [4], and ${\mathbf{u}}\in\mathcal{U}\subset\mathbb{R}^{M}$ is the design variable vector. Figure 1 illustrates the general concept of shape modification, showing how an original (or parent) geometry ${\mathbf{g}}({\boldsymbol{\xi}})$ defined by curvilinear coordinates ${\boldsymbol{\xi}}$, can be transformed (mapped) into a modified geometry ${\mathbf{g}}^{\prime}({\boldsymbol{\xi}},{\mathbf{u}})$ through a shape modification vector ${\boldsymbol{\delta}}({\boldsymbol{\xi}},{\mathbf{u}})$. Specifically, this vector captures the geometric displacement (modification) of each point on the surface (or manifold) due to changes in the design variables ${\mathbf{u}}$. This example highlights the flexibility and explicit nature of the shape optimization procedure, clarifying visually the geometric implications of manipulating the parameter vector ${\mathbf{u}}$.
Figure 2: Typical simulation-based design optimization block diagram for shape optimization
@@
$\min_{{\mathbf{u}}}\quad$ $f({\mathbf{u}})$ (2)
subject to $g_{i}({\mathbf{u}})\leq 0,\quad i=1,\ldots,I$
and to $h_{j}({\mathbf{u}})=0,\quad j=1,\ldots,J$
@@
where $f$ is the objective function, ${\mathbf{u}}_{l}$ and ${\mathbf{u}}_{u}$ are the lower and upper bounds for the design variable vector, $g_{i}$ the inequality constraints, and $h_{j}$ the equality constraints. Figure 2 provides a high-level flowchart of a typical simulation-based design optimization procedure for shape optimization. The process starts with the initial definition of the design variables and constraints, followed by a loop of design generation (shape modification), numerical simulations (solver), evaluation of the objective and eventual constraint functions, and optimization (optimizer). This iterative cycle is essential to systematically refine the geometry and achieve optimal design solutions that fulfill specified performance criteria.
@@ 3 Tackling the Curse of Dimensionality
The curse of dimensionality, a term coined by Richard Bellman in 1961, describes the exponential increase in complexity that accompanies adding dimensions to a mathematical problem [1]. In the context of shape optimization, as the dimensionality $M$ of the design variable vector $\mathbf{u}$ (along with, to some extent, the design space bounds) increases, the volume of the space grows exponentially, making it increasingly difficult to cover it adequately with a finite number of observations. This complexity not only impacts the efficacy of optimization methods, see, e.g., Fig. 3, including those relying on surrogate models and multi-fidelity approaches but also complicates surrogate training and uncertainty quantification.
@@
Figure 4 provides a systematic classification of dimensionality reduction techniques applied in shape optimization. Addressing the curse of dimensionality in such a complex domain requires distinguishing between two principal strategies: space reduction and dimensionality reduction. Each branch clearly highlights the specific methods discussed in this survey, providing readers with a comprehensive framework to navigate the broad landscape of techniques explored in subsequent sections
@@
Furthermore, design-space dimensionality reduction techniques can be categorized based on their reliance on simulation or geometric data. Simulation-driven techniques, such as indirect methods and some forms of GTM, derive their effectiveness from detailed understanding and manipulation of the objective functions in simulations. Conversely, methods like PCA and its variants primarily utilize geometric data. However, modern approaches often blend these data types, leveraging both geometric and physical simulation data to enhance the applicability and effectiveness of the dimensionality reduction [11].
The choice of dimensionality reduction techniques in shape optimization significantly influences the efficiency and success of the optimization process. By classifying these methods into distinct categories and understanding their underlying principles and applications, engineers and researchers can better select and apply the most appropriate techniques for their specific design challenges. This classification not only aids in comprehending the vast landscape of dimensionality reduction but also highlights the innovative ways these techniques are integrated and applied in the complex field of shape optimization of functional surfaces.
It should be noted that the primary goal of this review is to systematically map which dimensionality reduction methods are currently employed within shape optimization, given the unique severity of the curse of dimensionality in this context due to expensive performance evaluations. While certain techniques, particularly those within the PCA family, show specific adaptations tailored for shape optimization problems, others are employed directly, with no significant modifications compared to their general usage in other engineering domains.
@@ 3.1 Literature overview
To empirically assess the prevalence and efficacy of design space dimensionality reduction methods, a scoping review [12, 13, 14] was conducted, focusing on their applications across various domains of shape optimization. Scoping reviews, arisen in clinical medicine, use a broad approach for mapping literature and addressing broad research questions without performing articles’ quality assessment [15].
3.1.1 Research Questions
The scoping review focuses on the examination of dimensionality reduction methods used in shape optimization, specifically aimed at addressing the challenges associated with high-dimensional design spaces. The following research questions guide this investigation:
  1. 1.
    Which dimensionality reduction methodologies are currently applied to manage the curse of dimensionality in shape optimization problems?
  2. 2.
    What advantages and limitations are associated with applying dimensionality reduction techniques to shape optimization?
  3. 3.
    What emerging directions and opportunities exist for future research and development in dimensionality reduction for shape optimization?
3.1.2 Inclusion and Exclusion Criteria
The inclusion criteria for this review were carefully defined to identify literature directly addressing shape optimization problems through dimensionality reduction methods. Conversely, the exclusion criteria aimed to ensure the relevance and quality of the selected articles by omitting studies not explicitly related to shape optimization. Additionally, non-peer-reviewed contributions, such as conference papers, abstracts, editorials, and other grey literature, were excluded to maintain academic rigor.
3.1.3 Databases and Keywords
Scopus and Web of Science (WoS) were selected as primary databases for this review due to their extensive coverage of interdisciplinary literature, thus providing a comprehensive and rigorous foundation for identifying relevant studies in shape optimization.
The bibliographic search strategy employed targeted combinations of specific keywords to effectively capture relevant research on dimensionality reduction techniques within shape optimization. The query terms used were: (‘‘shape optimi*ation’’ OR ‘‘form optimi*ation’’ OR ‘‘design optimi*ation’’ OR ‘‘geometr* optimi*ation’’) AND (‘‘dimension* reduc*’’ OR ‘‘space reduc*’’ OR ‘‘reduce* dimension*’’ OR ‘‘reduced number of dimens*’’ OR ‘‘reduced design variable*’’ OR ‘‘reduced number of design variable*’’). These keywords were strategically selected and searched within article titles, abstracts, and keywords sections (using TITLE-ABS-KEY filter for Scopus and TS filter for WoS), ensuring the capture of relevant articles specifically addressing dimensionality reduction methods in shape optimization.
Figure 5: PRISMA flow chart
3.1.4 Search Procedure
The search procedure followed the preferred reporting items for systematic reviews and meta-analyses extension for scoping reviews (PRISMA-ScR) guidelines [16]. The process is depicted in the PRISMA-ScR flow diagram (see Fig. 5), detailing the steps from initial identification to the final selection of articles.
Initially, a total of 435 records were identified through database searches conducted on Scopus and Web of Science (WoS). To further enhance comprehensiveness, reference lists of the included articles were manually examined to identify additional relevant studies.
Following the initial search, a screening phase assessed the relevance of titles and abstracts, resulting in the exclusion of 278 records and retaining 157 potentially relevant articles for detailed review. Subsequently, during the eligibility assessment phase, a full-text analysis was performed, applying specific inclusion and exclusion criteria. Articles were excluded if they did not explicitly address dimensionality or space reduction methods applied to shape optimization problems.
Ultimately, 112 articles were selected, representing a rigorously curated set of literature directly relevant to the focus of this review.
@@ 3.1.4 Search Procedure
3.2 Scoping Review Outcomes
The statistical analysis, supported by visual trends from the literature, reveals distinct patterns in the adoption of direct versus indirect methods, the preference for linear over nonlinear approaches, and the allocation between space reduction and dimensionality reduction techniques. This analysis not only reflects the current state-of-the-art but also guides future research directions by identifying areas lacking in-depth exploration and potential methodological synergies.
@@ 3.2 Scoping Review Outcomes
When distinguishing between space reduction and dimensionality reduction, as shown in Fig. 7, 79% of the studies focus on dimensionality reduction while only 21% address space reduction. This highlights a predominant trend towards techniques that reduce the number of variables directly, rather than just condensing the space within which the variables exist.
Considering dimensionality reduction methods, the examination of literature trends underscores a significant leaning towards direct dimensionality reduction methods, with a staggering 93% of reviewed articles opting for direct techniques as opposed to only 7% utilizing indirect methods. This suggests a strong preference for methods that actively transform the input data space to achieve dimensionality reduction (see Fig. 7).
In terms of methodological choices for direct dimensionality reduction approaches (see Fig. 9), there is an overwhelming preference for linear techniques, with 80% of the cases applying linear methods over 20% for nonlinear approaches. This could indicate that the simplicity and computational efficiency of linear methods, such as PCA and SVD, continue to make them popular choices in many engineering applications despite the increasing capability of nonlinear methods to capture more complex data relationships.
Finally, analyzing the approaches based on their reliance on data types, as shown in Fig. 9, 49% of the methods are predominantly geometry-based, 38% are simulation-based, and a smaller segment of 13% combines both approaches. This distribution emphasizes the critical role of geometric considerations in the design processes and the significant integration of simulation data to inform these designs.
@@
The temporal trend across the years (see Fig. 10) shows a cubic trend in the increment of the publications related to dimensionality reduction in shape optimization (note that the data for the year 2025 is partial, as the bibliographic research was conducted on February 26, 2025). This growing interest underlines the expanding recognition of the importance of advanced dimensionality reduction techniques in tackling the complexity of modern engineering design problems.
@@ 4 Space Reduction Methods
Space reduction methods play a crucial role in the field of shape optimization, particularly when dealing with high-dimensional design spaces that challenge traditional optimization techniques. These methods enhance the efficiency and effectiveness of optimization processes by strategically narrowing the focus of the design space. This is achieved either through bounds narrowing, which reduces the breadth of the design space to concentrate on regions most likely to contain optimal solutions, or through geometric filtering, which excludes non-viable or abnormal regions based on sophisticated data-driven models or constraints. Figure 11 visually contrasts the two main categories of space reduction methods for a simplified two-dimensional design space ($u_{1},u_{2}$). Both approaches aim to refine the search area, ensuring that optimization efforts are concentrated on the most promising or feasible parts of the design space. By focusing on these key strategies, the following sections will explore how bounds narrowing and geometric filtering serve as effective tools in simplifying complex optimization tasks, ultimately leading to more innovative and practical engineering solutions.
@@ 4.1 Bounds Narrowing
These methods are instrumental in narrowing the bounds of traditionally expansive design spaces, focusing on smaller, more promising subspaces where optimal solutions are likely to be found.
In aerodynamic shape optimization, bounds narrowing plays a critical role, employing adaptive response surface methodologies and intelligent exploration strategies [17, 18, 19]. Kriging-based surrogate models, constructed and refined within reduced subspaces, efficiently concentrate on key performance regions, significantly streamlining the optimization process. A variable-fidelity optimization framework, coupled with an intelligent design space reduction strategy, has demonstrated substantial computational time savings while preserving high-fidelity accuracy in airfoil optimization problems [20]. Similarly, a variable-accuracy global metamodel-based optimization approach has been proposed, effectively narrowing the bounds of design experiments and uncertainty quantification for fluid-structure interaction problems [21].
In multidisciplinary optimization contexts, bounds narrowing methods effectively reduce complexity and computational cost. For instance, in optimizing an air-launched satellite vehicle, genetic algorithms combined with simulated annealing successfully narrowed the search space to promising multidisciplinary parameter regions [22]. Similar methodologies have also been applied to wing-weight minimization and aero-structural decomposition problems, significantly enhancing computational efficiency [23, 24].
In naval and underwater vehicle design, bounds narrowing has been effectively implemented using set-based and hierarchical space reduction methods. Set-based design approaches have been employed in ship design, maintaining flexibility and adapting to evolving requirements within systematically reduced design spaces [25]. Dynamic [26] and hierarchical [27, 28] space reduction frameworks have further advanced ship hull optimization, utilizing data mining techniques such as partial correlation analysis [29, 30], self-organizing maps [28], and rough set theory [31] to considerably improve optimization efficiency and accuracy. Innovative hierarchical and ensemble surrogate-based optimization methods have significantly improved optimization performance in hydrodynamic shape optimization of blended-wing-body underwater gliders [27, 32]. Furthermore, bounds narrowing has effectively guided optimization efforts toward critical design regions, demonstrating notable hydrodynamic performance gains for underwater gliders [33].
In structural optimization, a novel method integrating particle swarm optimization with chaotic logistic maps and comprehensive learning strategies has refined dynamically reduced spaces, effectively enhancing the ability to escape local optima and improving the search process for practical-scale engineering problems [34].